OpenStudy (anonymous):

Solve this differential equation: y''+9y=15.25cos(3.1x)

OpenStudy (anonymous):

First solve the homogenous problem. Then find a particular solution to this.

OpenStudy (anonymous):

\[ y \left( x \right) =C_1 \sin \left( 3\,x \right)+C_2\cos \left( 3 \,x \right) -25\,\cos \left( {\frac {31}{10}}\,x \right) \]

OpenStudy (anonymous):

yes i understand that is the solution but can you show me how you arrived at it because i also solved it in wolfram

OpenStudy (anonymous):

can you show me how we get the particular solution because i already have the complementary solution